OpenStudy (anonymous):

calculus help...plz explain and answer question...i totally dont know what they are...studying on my own

OpenStudy (fibonaccichick666):

OOOH yay! ok lol so do you know how to parametrically define a vector?

OpenStudy (anonymous):

not really lol

OpenStudy (fibonaccichick666):

ok lol give me a minute to write something up

OpenStudy (anonymous):

sure lol

OpenStudy (anonymous):

would you just find those things out....btw you have a sweet voice

OpenStudy (fibonaccichick666):

yep you just find \[\frac{d}{dt}\] for both the x and y coordinate

OpenStudy (anonymous):

1-cost 1+sint

OpenStudy (fibonaccichick666):

what was the 1st deriv of the y again?

OpenStudy (fibonaccichick666):

hint redo it

OpenStudy (anonymous):

i think the one i gave u is are first derivatives

OpenStudy (fibonaccichick666):

the first deriv of y is wrong

OpenStudy (anonymous):

second derivative would be sint cost

OpenStudy (anonymous):

i m confused srry lol

OpenStudy (fibonaccichick666):

ughh this is so hard to do on here!!! ok so pardon if i screw it up \[<{t-sin~t},{1-cos~t}>\] This is your position vector. Now we want the velocity vector so:\[ \color{blue}{<\frac{d}{dt}(t-sin~t),\frac{d}{dt}(1-cos~t)>} \] Then acceleration is the next derivative

OpenStudy (anonymous):

hey to be honest can u do it on educreations...ur a great teacher and ur voice is so sweet ... plzzz

OpenStudy (anonymous):

<1-cos(t),sin(t)> - first derivative <sin(t),cos(t)> - second derivative

OpenStudy (fibonaccichick666):

haha gtg maybe later

OpenStudy (anonymous):

ding it...XD...kk bye